OpenStudy (anonymous):

If AB, across the window, measures 25 inches, calculate the area of the window in square inches.

OpenStudy (anonymous):

Now when we apply Pythagoras. Lets say side is a.\[ \sqrt{a^2 + a^2 } = 25\]\[\sqrt2a=25\]\[a = \frac{25}{\sqrt2}\]

OpenStudy (anonymous):

Ares of square is \(a^2\) Area is \(\frac{625}{2}\)

OpenStudy (anonymous):

so wats the answer...im not getting this

OpenStudy (anonymous):

at all

OpenStudy (lalaly):

is ther a picture?

OpenStudy (anonymous):

Area is \(\frac{625}{2}\)

OpenStudy (lalaly):

whats the shape of the window?

OpenStudy (anonymous):

its just a circle in the pic...

OpenStudy (lalaly):

A = pi r^2 25^2pi = 625 625pi

OpenStudy (anonymous):

Okay I thought A square sorry. Lana is correct

OpenStudy (anonymous):

haha thanks...i still have a few questions